The length of a rectangular scarf is 9 inches more than five times its width. The area of the scarf is 308 square inches. Determine the length and width of the scarf?
Answer by stanbon(75887)

The length of a rectangular scarf is 9 inches more than five times its width. The area of the scarf is 308 square inches. Determine the length and width of the scarf?
------------
Let the width be "x" inches.
Length = "5x+9" inches
-------------
Area = width * length
Area = x * (5x+9) = 308
5x^2+9x-308=0
x=[-9+-sqrt(81-4*5*-308)]/10
x=[-9+-sqrt6241]/10
x=[-9+-79]/10
Positive Answer: x=7 inches (this is the width of the scarf)
5x+9 = 44 inches (this is the length of the scarf)
===========
Cheers,
Stan H.